# Schur’s Positive-Definite Network: Deep Learning in the SPD cone with structure

### v1

@@ 5 Experiments
All three models are trained using ADAM with default parameters (Kingma and Ba, 2014) to minimize a reconstruction error loss $\mathcal{L}_{\mathrm{MSE}}=\frac{1}{N_{\mathrm{train}}}\sum^{N_{\mathrm{train}}}_{i=1}\|\hat{\bm{\Theta}}^{(i)}-\bm{\Theta}^{(i)}_{\mathrm{true}}\|^{2}_{F}$ in the vein of GLAD, where $N_{\mathrm{train}}$ is the size of the training set and $\hat{\bm{\Theta}}$ is the model’s output.
Architectures & training
We consider in our experiments models with $K=1$ SpodNet layer. The input to the models are the empirical covariance matrices $\bm{S}$, from which we compute $\bm{\Theta}_{\mathrm{init}}=(\bm{S}+0.1*\bm{I})^{-1}$ and $\bm{W}_{\mathrm{init}}=\bm{S}+0.1*\bm{I}$, which serve as the inputs to the SpodNet layer. We use learning rates of $0.01$, $0.02$ and $0.03$ respectively, to train UBG, PNP and E2E. The principal difficulty in training UBG and PNP stems from learning to predict $\gamma^{+}$. We find that using a perceptron with a single hidden layer of 10 neurons and ReLU activations that takes in the current state of the whole matrix to predict a $\gamma^{+}$ is expressive enough to learn adequate matrices (Figures 4 and 6). Its positivity is ensured by taking the absolute value of the prediction. We believe that a more extensive architectural search for this perceptron could help achieve even better results. The prediction of the $\Lambda^{+}$ parameters at each step is performed by a small single layer perceptron of 5 neurons with ReLU activation functions and an absolute value function following the last layer to ensure positivity. This network acts elementwise on the entries of the input vector. The sufficiency of a lower amount of training data for UBG could be imputed to its stronger inductive bias. PNP’s $\Psi$ function and E2E’s $\Phi$ function are learned by a single layer ReLU perceptron of 500 neurons. All experiments were ran on an RTX 2080 SUPER GPU with an AMD EPYC 7302 CPU and resources of similar computing power.
Sparsity & positive-definiteness
Figure 5 highlights that all three of our models are indeed able to learn jointly sparse and SPD matrices. Contrary to GLAD, the raw predictions respect these structural constraints by construction. For the sake of completeness, we threshold the entries of GLAD’s raw predictions at increasing values to artificially introduce sparsity, and observe that this can indeed break their positive-definiteness. As visual confirmation, Figure 6 shows one of UBG and PNP’s predictions in the weakly sparse and strongly sparse settings.
Performance & benchmark
Since GLAD’s predictions are not naturally sparse and SPD, we benchmark against sklearn’s implementation of the GLasso solver (Friedman et al., 2008), which is the standard algorithm for learning sparse precision matrices. We compare ourselves to an ideal Oracle GLasso where the $\ell_{1}$ hyperparameter $\lambda^{(i)}$ is individually assigned to the solver for each target matrix $\bm{\Theta}_{\mathrm{true}}^{(i)}\forall i\in[N_{\mathrm{test}}]$ and where $\lambda^{(i)}$ is optimally tuned so as to achieve best possible reconstruction error. In practice, we validate this $\lambda^{(i)}$ in $[10^{-3}\lambda_{\max},\lambda_{\max}]$ where $\lambda_{\max}$ is the parameter such that every off-diagonal entry of the prediction is zero (see Section A.2). We emphasize that this is an idealistic scenario where the GLasso is at its most favorable. The results are depicted in Figure 4 for the weakly sparse setting and illustrates that our three models are indeed able to learn jointly sparse and SPD matrices while outperforming the Oracle GLasso solver. PNP yields the best performance in terms of NMSE whereas the E2E model is best in terms of support recovery in our setting. Figure 6 offers visual confirmation of the relevancy of the predictions. Interestingly, merely only one SpodNet layer is enough to achieve good performances for all of the three models.
Figure 5: Sparsity (left) and positive-definiteness (right) of the raw predictions of all three of our models and GLAD’s, along with the latter’s thresholded predictions, on the testing set of the weakly sparse setting.
Figure 6: Comparison of sparsity degrees and smallest eigenvalues of UBG and PNP’s predictions with an Oracle GLasso, in both the strongly sparse (left panel) and weakly sparse (right) settings.
6 Conclusion & future directions
@@
We have proposed Schur’s Positive-Definite Network (SpodNet), a novel learning module compatible with other standard architectures offering strict guarantees of SPD outputs. The principal novelty of SpodNet comes from its higher expressivity than existing approaches, as it is also able to handle additional desirable structural constraints. In this work, we highlighted that it can be used to efficiently learn sparse and SPD matrices. To the best of our knowledge, SpodNet layers are the first to offer strict guarantees of such highly non-trivial structure in the outputs. We have shown how to leverage these guarantees to build several neural architectures that outperform the standard solvers in the context of sparse precision matrix estimation. Current directions of research include algorithmic improvements in order to speed up SpodNet’s column/row/diagonal update framework and theoretical investigations in order to better understand the spectral perturbations undergone by the inputs.
@@ 7 Acknowledgments
We would like to thank Badr Moufad (Ecole Polytechnique, France) for his help with the PyTorch implementations of SpodNet, Paulo Gonçalves (Inria Lyon, France) for fruitful discussions as well as the Centre Blaise Pascal for computing ressources, which uses the SIDUS solution developed by Emmanuel Quemener (ENS Lyon, France) (Quemener and Corvellec, 2013). This work was partially supported by the AllegroAssai ANR-19-CHIA-0009 project.
Appendix A Supplementary Material
A.1 GLAD’s update rules
Each unrolled iteration of GLAD, which can be seen as an individual layer in a deep learning perspective, updates several running variables:
@@ A.1 GLAD’s update rules
A.2 Critical regularization parameter for the Graphical Lasso
Proposition A.1.
Let $\hat{\bm{\Theta}}=\operatorname*{argmin}_{\bm{\Theta}}-\log\det\bm{\Theta}+\langle\bm{S},\bm{\Theta}\rangle+\lambda\|\bm{\Theta}\|_{1,\mathrm{off}}$. Then, $\hat{\bm{\Theta}}$ is diagonal if and only if $\lambda\geq\lambda_{\max}=\max_{i\neq j}|s_{ij}|$.
Proof.
The first-order optimality condition, aka Fermat’s rule, for the Graphical Lasso is:
$\hat{\bm{\Theta}}=\operatorname*{argmin}_{\bm{\Theta}}-\log\det\bm{\Theta}+\langle\bm{S},\bm{\Theta}\rangle+\lambda\|\bm{\Theta}\|_{1,\mathrm{off}}\Leftrightarrow\hat{\bm{\Theta}}^{-1}+\bm{S}\in\partial\|\hat{\bm{\Theta}}\|_{1,\mathrm{off}}$ (19)
Now $\bm{U}\in\partial\|\hat{\bm{\Theta}}\|_{1,\mathrm{off}}$ if and only if
  • $U_{ii}=0$ for $i\in[p]$
  • $U_{ij}\in\lambda\partial|\Theta_{ij}|=\begin{cases}\lambda\operatorname{sign}(\Theta_{ij})$ if $\Theta_{ij}\neq 0\\ [-\lambda,\lambda]\text{ otherwise}\end{cases}$ for $i,j\in[p],i\neq j$.
Hence
[…]
$\Leftrightarrow s_{ij}\in[-\lambda,\lambda]\quad\forall i\neq j$ (21)
$\Leftrightarrow\lambda\geq\lambda_{\max}=\max_{i\neq j}|s_{ij}|$ (22)

### v3

@@ 5.1 Sparse precision matrix recovery
We train our three models (UBG, PNP and E2E) to minimize a reconstruction error $\mathcal{L}_{\mathrm{MSE}}=\frac{1}{N_{\mathrm{train}}}\sum^{N_{\mathrm{train}}}_{i=1}\|\hat{\bm{\Theta}}^{(i)}-\bm{\Theta}^{(i)}_{\mathrm{true}}\|^{2}_{F}$ in the vein of Gregor and LeCun (2010); Shrivastava et al. (2020), where $\hat{\bm{\Theta}}$ is the model’s output. All three models are trained using ADAM with default hyperparameters (Kingma and Ba, 2014).
Baselines
We compare our three models to various other approaches throughout our experiments: learning-based, Riemannian and model-based estimators. As performed in Belilovsky et al. (2017); Shrivastava et al. (2020), we use the three most popular methods for estimating precision matrices as our core baselines: the GLasso (Friedman et al., 2008), the inverse Ledoit-Wolf estimator (Ledoit and Wolf, 2004) and the inverse OAS (Chen et al., 2010). The Ledoit-Wolf and OAS estimators are covariance matrix estimators, for which we take the inverse for estimating the precision (we simply refer to them as Ledoit-Wolf and OAS). All three approaches guarantee the positive-definiteness of the estimated matrix but the GLasso is the only one to additionally ensure sparse predictions. For the GLasso, for each matrix $\bm{S}^{(i)}$ a sparsity-regulating hyperparameter $\lambda^{(i)}$ is chosen using sklearn’s GraphicalLassoCV cross-validation procedure, that relies only on the samples used in $\bm{S}^{(i)}$. In the same vein, we use dedicated Ledoit-Wolf and OAS estimators for each $\bm{S}^{(i)}$. We also compare our approaches to the GLAD model (Shrivastava et al., 2020), considering both its $\bm{\Theta}$ and $\bm{Z}$ outputs. We recall that $\bm{Z}$ is not guaranteed to be positive-definite and its validity as a precision matrix estimator is therefore debatable. GLAD is trained with the same loss, data and optimizer as our models.
We compare the performance of the different approaches in terms of Normalized MSE on the test set: $\mathrm{NMSE}=\frac{1}{N_{\mathrm{test}}}\sum^{N_{\mathrm{test}}}_{i=1}\frac{\|\hat{\bm{\Theta}}^{(i)}-\bm{\Theta}^{(i)}_{\mathrm{true}}\|^{2}_{F}}{\|\bm{\Theta}^{(i)}_{\mathrm{true}}\|^{2}_{F}}$ and F1 score for assessing support recovery. For the latter, false positives correspond to $\hat{\bm{\Theta}}_{ij}\neq 0$, $(\bm{\Theta}_{\mathrm{true}})_{ij}=0$ (and similarly for true positives and false negatives).
As shown in Figure 4, at convergence, the learning-based methods show to be more suited for the actual reconstruction of $\bm{\Theta}_{\mathrm{true}}$ as they significantly outperform traditional methods. This superior performance is especially pronounced as the dimensionality of the data grows and the sparsity of the ground truth decreases. In the most challenging scenario ($p=100$ and $n=100$ with weakly sparse $\bm{\Theta}_{\mathrm{true}}$), the traditional methods perform poorly, with NMSE around 90%.
Figure 4: Learning-based (in variations of blue) vs traditional methods (in variations of red). Dotted curves indicate when one of the constraints (SPDness or sparsity) is not guaranteed. First row: Strongly sparse $\bm{\Theta}_{\mathrm{true}}$. Second row: Weakly sparse $\bm{\Theta}_{\mathrm{true}}$.
Figure 5: Comparing models in sample deficient regimes ($n\leq p$) up to large-sample regimes ($n\gg p$), evaluated in terms of NMSE and F1 score for support recovery. Dotted curves indicate when one of the constraints (SPDness or sparsity) is not guaranteed. In large dimension, GLAD’s $\bm{Z}$ performs well, but is never SPD; GLAD’s $\bm{\Theta}$ is SPD, but performs badly.
Figure 4 also suggests that our models perform especially well in terms of NMSE in the low-samples regime ($n\leq p$), which is a common setting in various real-world applications, such as neuroscience time-series analysis or financial temporal network inference. In Figure 5, we thus study the influence of the number of samples $n$ on the performance of each method. The results show that although all methods achieve excellent recovery in the large sample regime (down to around 2.5% NMSE in the lower dimensional $p=20$ setting), our models are especially more suited for matrix reconstruction in case of sample-deficiency, especially so in the higher dimensional ($p=100$) setting. In terms of support recovery, our models achieve significantly better F1 scores than the GLasso as the number of sample $n$ grows, with up to 25% improvement in certain tested settings. In the lower dimensional setting ($p=20$), we are even able to achieve a F1 score of over 0.75 with our UBG model. Finally, although GLAD’s $\bm{Z}$ performs well, an additional experiment in Section B.3 shows that those matrices turn out to never be SPD in various settings, making them unsuited as precision matrix estimators.
5.2 Application to unsupervised graph learning on a real-world dataset
We finally evaluate the performance of SpodNet in a graph learning context using a real-world dataset and under an unsupervised scenario. The aim of this experiment is twofold: (1) to assess the generalization capabilities of SpodNet, and (2) to determine if it can yield an accurate graph topology when trained solely on synthetic data, following the idea of Belilovsky et al. (2017).
Figure 6: Graphs estimation on the Animals dataset with the GLasso, EGFM and our UBG model. The former and the latter yield cleaner and more interpretable clusters than the GLasso, but EGFM is specifically tailored for this task as opposed to our UBG model.
Data & training
We consider the Animals dataset (Lake and Tenenbaum, 2010), which comprises $p=33$ animal species, each characterized by responses to $n=102$ binary questions (e.g., "Has teeth?", "Is poisonous?"). The objective is to infer a graph of connections between these $p$ animals based on the $n$ responses. We train our UBG model by minimizing the MSE on the synthetic data detailed in Section 4.2, with $p=33,n=102$ and various levels of sparsity. We generate 1000 training matrices and train the model for 100 epochs. The predicted precision matrix is then computed by inputting the empirical covariance matrix from the Animals dataset into SpodNet.
Baselines
We compare our approach with the GLasso (with the regularization parameter selected by cross-validation), and the Elliptical Graphical Factor Model (EGFM) proposed by Hippert-Ferrer et al. (2023). The latter is a Riemannian optimization method specifically designed to identify clusters within the data by estimating a precision matrix whose inverse is modeled as a low-rank matrix plus a positive diagonal. For a fair comparison, we adopt the same hyperparameters as outlined in the original paper, setting the rank and regularization parameter to 10.
Each method yields a precision matrix, which is used to construct a graph representing the connections between the $p$ animals. The graph is constructed by considering the absolute values of the precision matrix coefficients, which represent the strength of the connections, and by removing the diagonal elements to eliminate self-loops. The results are illustrated in Figure 6, where the nodes of the graphs are partitioned using the Louvain algorithm (Blondel et al., 2008). Qualitatively, we observe that UBG produces a coherent graph, with outcomes comparable to those of GLasso and EGFM. Additionally, the graph learned by our UBG model exhibits a cleaner structure compared to GLasso, with a sparser representation and finer clusters, such as the grouping of (Chimp, Gorilla). To quantitatively assess the quality of the obtained graphs, we calculate the modularity $m$ of the partitions: higher modularity values indicate better separation of the graph into distinct subcomponents (Newman, 2006). We find that the graph produced by UBG achieves a higher modularity ($m=0.78$) than GLasso ($m=0.61$) and slightly lower than EGFM ($m=0.86$). It is important to note that the EGFM is specifically designed to obtain clustered graphs, whereas our UBG model has no such inductive bias. These combined quantitative and qualitative results demonstrate that SpodNet generalizes effectively to unseen data and produces a coherent graph structure.
6 Conclusion
@@
We have proposed Schur’s Positive-Definite Network (SpodNet), a novel learning module compatible with other standard architectures, offering strict guarantees of SPD outputs. The principal novelty of SpodNet comes from its ability to handle additional desirable structural constraints, such as elementwise sparsity which we used as an illutrative example throughout this paper. To the best of our knowledge, SpodNet layers are the first to offer strict guarantees of such highly non-trivial structure in the outputs. In future works, SpodNet could be leveraged to learn other additional structures beyond sparsity. We have shown how to leverage SpodNet to build neural architectures that outperform traditional methods in the context of sparse precision matrix estimation in various settings. Future directions of research include theoretical investigations in order to better understand the expressivity and derive convergence guarantees for our approach.
7 Reproducibility statement
We believe this paper fully discloses the information needed to reproduce the main experimental results. Our core contribution is the SpodNet layer: its general algorithmic framework is presented in Algorithm 1. Three specific models built using this framework for learning sparse and SPD matrices are introduced in Section 4.2, and the exact architectures used for each of them throughout our experiments are explained in detail in Section A.2. The conducted experiments are described in Section 5 with additional details in Appendix A. PyTorch implementations are provided along with the submission.
@@ 8 Acknowledgments
We would like to thank Badr Moufad (Ecole Polytechnique, France) for his help with the implementations, Paulo Gonçalves (Inria Lyon, France) for fruitful discussions as well as the Centre Blaise Pascal for computing ressources, which uses the SIDUS solution developed by Emmanuel Quemener (ENS Lyon, France) (Quemener and Corvellec, 2013). This work was partially supported by the AllegroAssai ANR-19-CHIA-0009 project.
Appendix A Experimental details
A.1 Synthetic data generation
Here we provide additional details on the experiment of Section 5.1. The sparsity degree of each matrix is controlled through the one imposed on their Cholesky factors during their generation, with the alpha parameter of the function. Throughout our experiments, we consider alpha $\in\{0.7,0.95\}$; this is the probability of a 00 entry on the Cholesky factor of the generated matrix and not the actual fraction of null elements of $\bm{\Theta}$. Although the numbers vary with the dimension $p$, $\alpha=0.95$ roughly leads to 90 % of zero entries, while $\alpha=0.7$ leads to 25 % of zero entries.
A.2 Architectures for UBG, PNP and E2E
The function $g$ for all three models is a MLP with two hidden layers of 3 neurons each with a ReLU activation function. It takes as input $\theta_{22},s_{22}$ and $(\bm{\theta}_{12}^{+})^{\top}[\bm{\Theta}_{11}]^{-1}\bm{\theta}_{12}^{+}$.
UBG
Its $\gamma^{+}$ parameter is predicted by a MLP that takes as input the current state of the $\bm{\theta}_{12}$ block (of dimension $p-1$) and has a single hidden layer of $\lfloor p/2\rfloor$ neurons with a ReLU activation function. The (single) output neuron has an absolute value activation, in order to keep it positive since it predicts a step-size.
UBG involves another MLP to learn the vector $\bm{\lambda}^{+}$ of elementwise soft-thresholding parameters. This MLP has a single hidden layer of 5 neurons, with ReLU activation function. For the same reason as before, the output layer also has an absolute value activation.
PNP
First, the step-size $\gamma^{+}$ is predicted by a MLP with the exact same architecture as the one in UBG. The learned operator $\Psi$ takes as input $\bm{\theta}_{12}-\gamma^{+}(\bm{s}_{12}-\bm{w}_{12})$, passes it through a single hidden layer of $2p$ neurons, followed by a ReLU activation function, and projects it back into a $p-1$ dimensional vector.
The architecture to predict $\bm{\lambda}$ is the same as for UBG. Our experiments show that multiplying its output by $0.1$ helps avoid local minimas and improves convergence.
E2E
The neural network $\Phi$ is a MLP that takes as input $\bm{\theta}_{12}$, passes it through a single hidden layer of $10p$ neurons follow by a ReLU activation function, and projects it back into a $p-1$ dimensional vector. The architecture to predict $\bm{\lambda}$ is the same as for UBG and PNP. Our experiments show that multiplying its output by $0.1$ helps avoiding local minimas and improves convergence.
GLAD
We use the authors’ original implementation retrieved from the authors repository at https://github.com/Harshs27/GLAD, with the default hyperparameters.
All four models are implemented with $K=1$ layer throughout all the experiments, since our results show that this was enough to yield competitive to outperforming results in the settings under consideration.
A.3 Details on Figures 4 and 5
For the weakly sparse settings, we use a learning rate of $10^{-3}$ for all three of our own models. For the strongly sparse settings, we use a learning rate of $10^{-2}$. GLAD’s learning rate is set to $10^{-2}$ in all settings. All four models are trained on the same 1000 training matrices, using a batch-size of 10, with ADAM’s default parameters. The models are tested on the same 100 matrices as mentioned in Section 5.1.
For Figure 5 We use the strongly sparse ground truths in order for the F1 score to be more sensical, and to better illustrate the relevancy of the learned sparsity patterns. We use a learning rate of $10^{-2}$ for all three of them in every setting for $p=100$, and of $5\cdot 10^{-3},10^{-2},10^{-3},3\cdot 10^{-2}$ in the $p=20$ setting for respectively $n=10,n=20,n=100,n=200$ and $n=500$.
Appendix B Additional details on models
B.1 SpodNet’s potential instabilities without normalization
We show in Figure 7 the evolution of the conditioning of an example of $\bm{\Theta}$ during training that becomes unstable if we do not use the stabilization procedure described in Section 4. We observe that although the matrix remains SPD as predicted by Proposition 3.1 but gets increasingly closer to being singular during the updates.
Figure 7: Potential instabilities without normalization. Along column update indices, we plot the smallest eigenvalue (blue), largest diagonal value (orange) and conditioning (green) of a $\bm{\Theta}$.
B.2 GLAD’s update rules
GLAD unrolls an algorithm that solves the following optimization problem:
Each iteration of GLAD, which can be seen as an individual layer in a deep learning perspective, updates several running variables:
@@ B.2 GLAD’s update rules
B.3 GLAD’s non-SPD outputs
As Figure 8 shows, in the setting of Figure 4, GLAD’s best performing outputs $\bm{Z}$ are SPD in virtually $0\%$ if the case strongly sparse case with $p=100$.
Figure 8: Percentage of SPD outputs among GLAD’s $\bm{Z}$ and $\bm{\Theta}$ outputs at convergence. As detailed in Section 4.1, $\bm{Z}$ is not guaranteed to be sparse.